can you solve this:
1/2 log2 x squared = log4 4x

Answer by lwsshak3(11628)


1/2 log2 x squared = log4 4x
..
change to base 10
(1/2)logx^2/log2=log4x/log4
(1/2)2logx/log2=(log4+logx)/log4
logx/log2=(log4+logx)/log4
LCD=log2log4
log4logx=log2(log4+logx)
log4logx=log2log4+log2logx
log4logx-log2logx=log2log4
logx(log4-log2)=log2log4
logx=log2log4/(log4-log2)
logx=.602
x=10^.602=3.999
or rounded,x=4
Check:with x=4
1/2 log x^2/log2
1/2 log 16/log2=2
log4x/log4
log16/log4=2
left-hand side = right-hand side